# EventHallusion: Diagnosing Event Hallucinations in Video LLMs

### v4

@@ 5.2 TCD Evaluation
5.2.2 Implementation Details.
5.2.3 Effectiveness of TCD
We apply our proposed TCD method on three prevalent VideoLLMs, namely LLaVA-NeXT [22], VideoChat2 [35] and VILA [29], and evaluate their performances on EventHallusion. As shown in Tab.2, equipping with our TCD, three VideoLLMs achieve consistent performance boost. This demonstrates that our TCD method can be seamlessly integrated with versatile VideoLLMs and reduce their event-related hallucinations without bells and whistles.
5.2.4 Explorations on Contrastive Sample Constructions
We further explore more alternative strategies to construct the contrastive counterpart in our TCD method. Recalling that our TCD aims at ambiguating the temporal cues of the original video, we further explore the effects of blurring the spatial cues by adding gaussian noises to video frames when constructing the contrastive counterpart. As shown in Tab.3, introducing Gaussian noises on top of TCD can further boost the overall performances. However, the extent of introducing Gaussian noises should be delibrately adjusted to prevent degradation in practice. Therefore, we do not choose to apply this strategy in our TCD method.
5.2.5 Ablations on Hyper-parameters of TCD
@@
Figure 8: Qualitative Results of the TCD method on our EventHallusion Benchmark.
5.2.6 Effect of Sampled Frames in Distorted Video.
We further explore the effects of the temporal downsampling ratio when constructing the contrastive video in our TCD. Specifically, we adopt LLaVA-NeXT-Video[22] as the backbone model with 32 frames sampled for the original video input and downsample the number of sampled frames to construct the contrastive video. This experiment is conducted on the binary classification task of EventHallusion. From Tab.4, we have the following observations: (1) Extreme temporal downsampling (i.e, 1 frame) can lead to severe performance degradation of our TCD approach. This phenomenon may arise because single-frame input significantly diminishes spatial cues, leading to distorted prediction logits during contrastive learning. (2) A low temporal downsampling rate (e.g., 16 frames) performs worse than higher rates (e.g., 4 and 8 frames). This indicates that narrowing the gap between the contrastive video and the original one weakens the advantages of our TCD approach.
5.3 Qualitative Results
We also present qualitative results comparing the baseline method with and without our TCD approach on the EventHallusion benchmark. As shown in Fig.8, The baseline method fails to generate accurate responses, whereas incorporating our TCD approach yields correct answers. Due to the page limitation, we provide more examples in our supplementary materials.
6 Conclusion
@@
Supplementary Material
1 Comparison with other video hallcination benmarks.
We exhibit the comparison between HallusionBench, VideoHallucer, and our EventHallusion in Figure 9. As shown in Figure 9, compared to HallusionBench and Videohallucer emphasize assessing the models’ sensitivity to events’ chronological orders, our proposed EventHallusion focuses on analyzing the event-related language prior and vision-language correlation bias encapsulated in existing VideoLLMs.
2 Automatic Evaluation For Detailed Description Matching
We apply GPT-4o to evaluate the detailed description matching task of our Entire, Mix, and Misleading categories respectively. For Entire, we prompt GPT-4o to judge if VideoLLMs’ detailed description of the video is consistent with the ground-truth rare event. And for Mix, we prompt GPT4o to focus on the occurrence of the rare-common mixed event. Finally, for Misleading, we prompt GPT4o to assess the consistency of the ground-truth event and search for any unrelated events. In addition, we set up the default system prompt to better guide GPT-4o’s evaluation, which is shown in Figure 10. The detailed prompts are shown in Figure  11, 12, 13, respectively.
@@ 2 Automatic Evaluation For Detailed Description Matching
3 Qualitative Examples
3.1 Data examples of EventHallusion.
We present examples of Entire, Mix, and Misleading in Figure 14, 15, 16, incorporating responses of VideoLLMs. As shown in these examples, VideoLLMs are prone to generate hallucinated answers under all of our settings, indicating their deficiency in addressing event hallucinations.
3.2 Examples of GPT-4o’s Evaluation.
We present examples of GPT-4o’s evaluation of detailed description matching task on Entire, Mix, and Misleading respectively in Figure 17, 18, 19. As shown in these examples, GPT-4o have the capacity to judge the accuracy of video events.
3.3 Examples of TCD.
We present examples of our TCD in mitigating video event hallucinations. Specifically, we exhibit the generated response with and without TCD for LLaVA-NeXT-Video and VideoChat2 on our EventHallusion in Figure 20 and 21.

### v5

@@ Introduction
Implementation Details.
Effectiveness of TCD
We apply our proposed TCD method on three prevalent VideoLLMs, namely LLaVA-NeXT (Li et al. 2024a), VideoChat2 (Li et al. 2024b), and VILA (Lin et al. 2024), and evaluate their performances on EventHallusion. As shown in Tab.2, equipping with TCD, three VideoLLMs achieve a consistent performance boost. This demonstrates that TCD can be seamlessly integrated with versatile VideoLLMs and reduce their event-related hallucinations without bells and whistles.
Explorations on Contrastive Sample Constructions
We further explore more alternative strategies to construct the contrastive counterpart in our TCD method. Recalling that our TCD aims at ambiguating the temporal cues of the original video, we further explore the effects of blurring the spatial cues by adding gaussian noises to video frames when constructing the contrastive counterpart. As shown in Tab.3, introducing Gaussian noises on top of TCD can further boost the overall performance. However, the extent of introducing Gaussian noise should be deliberately adjusted to prevent degradation in practice. Therefore, we do not choose to apply this strategy in our TCD method.
@@
Ablations on Hyper-parameters of TCD
@@
Effect of Sampled Frames in Distorted Video.
We further explore the effects of the temporal downsampling ratio when constructing the contrastive video in our TCD. Specifically, we adopt LLaVA-NeXT-Video(Li et al. 2024a) as the backbone model with 32 frames sampled for the original video input and downsample the number of sampled frames to construct the contrastive video. This experiment is conducted on the binary classification task of EventHallusion. From Tab.4, we have the following observations: (1) Extreme temporal downsampling (i.e., 1 frame) can lead to severe performance degradation of our TCD approach. This phenomenon may arise because single-frame input significantly diminishes spatial cues, leading to distorted prediction logits during contrastive learning. (2) A low temporal downsampling rate (e.g., 16 frames) performs worse than higher rates (e.g., 4 and 8 frames). This indicates that narrowing the gap between the contrastive video and the original one weakens the advantages of our TCD approach.
Conclusion
@@
Appendix A Comparison with Other Video Hallucination Benchmarks.
We exhibit the comparison between HallusionBench, VideoHallucer, and our EventHallusion in Figure 8. As shown in Figure 8, compared to HallusionBench and VideoHallucer, which emphasize assessing the models’ sensitivity to events’ chronological orders, our proposed EventHallusion focuses on analyzing the event-related language prior and vision-language correlation bias encapsulated in existing VideoLLMs.
Appendix B Automatic Evaluation for Detailed Description Matching
We apply GPT-4o to evaluate the detailed description matching task of our Entire, Mix, and Misleading categories, respectively. For Entire, we prompt GPT-4o to judge if VideoLLMs’ detailed description of the video is consistent with the ground-truth rare event. And for Mix, we prompt GPT4o to focus on the occurrence of the rare-common mixed event. Finally, for Misleading, we prompt GPT4o to assess the consistency of the ground-truth event and search for any unrelated events. In addition, we set up the default system prompt to better guide GPT-4o’s evaluation, which is shown in Figure 9. The detailed prompts are shown in Figure  10, 11, 12, respectively.
@@ Appendix B Automatic Evaluation for Detailed Description Matching
Appendix C Qualitative Examples
Data examples of EventHallusion. We present examples of Entire, Mix, and Misleading in Figure 13, 14, 15, incorporating responses of VideoLLMs. As shown in these examples, VideoLLMs are prone to generate hallucinated answers under all of our settings, indicating their deficiency in addressing event hallucinations.
Examples of GPT-4o’s Evaluation. We present examples of GPT-4o’s evaluation of detailed description matching task on Entire, Mix, and Misleading respectively in Figure 16, 17, 18. As shown in these examples, GPT-4o has the capacity to judge the accuracy of video events.
Examples of TCD. We present examples of our TCD in mitigating video event hallucinations. Specifically, we exhibit the generated response with and without TCD for LLaVA-NeXT-Video and VideoChat2 on our EventHallusion in Figure 19, 20, and 21.